Classical Algebra Assignment

due Monday, January 29, 2001

Reading

Read section 3C in the textbook by Childs.

Exercises

1. 3C: 4

2. For each of the parts of 3C: 4 find the continued fraction expansion of the rational number a/b where a is the first listed integer and b the second.

3. Find the continued fraction expansion of the real number 2 − \sqrt{3} .

[Processed from GELLMU to L^{A}TEX on Fri 26 Jan 2001 at 09:16:39 EST]